Mr. Chai withdraws $18,000 from his saving account. Half of the money is given to his wife and the
rest is divided among his three children Andy, Ben and Charlie in the ratio 3:4:1 respectively.
(a) Ben saves 2/3 of his money. How much does he save?
(b) Charlie receives additional $1000 from his mother. Find the ratio of money received by Andy
to Charlie now.

Answer by mananth(16949)

Mr. Chai withdraws $18,000 from his saving account. Half of the money is given to his wife
wife gets $9000
rest is diczxfxcx]zfcfvided among his three children Andy, Ben and Charlie in the ratio 3:4:1 respectively.
Charlie gets 1/8 of 9000 = $1125
Ben gets 1/2 of 9000 = $4500
Andy gets 3/8 of 9000 = $3375
(a) Ben saves 2/3 of his money.
2/3 * 4500 = $3000
(b) Charlie receives additional $1000 from his mother.
Charlie gets 1125 +1000 = $2125
Andy gets $3375 Charlie gets $2125
Find the ratio of money received by Andy to Charlie now.
3375/1125
3:1